Find the equation of the line which contains the point (-2,3) and is parallel to the line 3x+5y=17. (write the numerical coefficient of each term to complete the required equation)
Answer would be: ___x + ___ y - ___ = 0

Found 2 solutions by josgarithmetic, Theo:
Answer by josgarithmetic(39617)
Answer by Theo(13342)

you want the line parallel to 3x + 5y = 17 that passes through the point (-2,3).

if the line is parallel, it will have the same slope.

convert 3x + 5y to slope intercept form of y = mx + b.

subtract 3x from both sides of the equation to get 5y = -3x + 17

divide both sides of the equation by 5 to get y = -3/5 * x + 17/5.

the slope is -3/5.

your parallel line must have the same slope.

your parallel line with have a slope of -3/5.

the point slope form of your equation will be (y - y1) = m * (x - x1).

your slope is equal to m which is equal to -3/5.
your point is equal to (-2,3) which is equal to (x1,y1).

when m = -3/5 and y1 is equal to 3 and x1 is equal to -2, the point slope form of the equation becomes:

(y - 3) = -3/5 * (x + 2)

multiply both sides of this equation by 5 to get:

5 * (y - 3) = -3 * (x + 2)

simplify to get:

5y - 15 = -3x - 6

add 3x to both sides of this equation and add 15 to both sides of this equation to get:

5y + 3x = 15 - 6

simplify and place x variable ahead of y variavble to get:

3x + 5y = 9

subtract 9 from both sides of this equation to get:

3x + 5y - 9 = 0

that's your solution.